# Intelligent pixel detectors: towards a radiation hard ASIC with on-chip machine learning in 28 nm CMOS

### v1

@@ 2 Single Cluster $p_{\mathrm{T}}$ Filtering Algorithm
Three models were developed with varying complexity to test how additional information improves $p_{\mathrm{T}}$ discrimination. The models were built to predict whether a cluster arose from three classes of charged particles: $p_{\mathrm{T}}>200$ MeV, and $p_{\mathrm{T}}<200$ MeV for positively and negatively charged particles. Two background classes were used because oppositely charged particles in a magnetic field were found to leave different cluster shapes. The threshold of 200 MeV was chosen to ensure flat signal efficiency for tracks with $p_{\mathrm{T}}>2$ GeV, a threshold important for physics analysis based on min-bias studies [7]. Model 1 is a two layer dense nueral network (DNN) with only 2 inputs , the cluster position ($y_{0}$) and size $y$-size. It correctly selects roughly 85% of tracks with $p_{T}>2$ GeV. Model 2 uses the full cluster $y$-profile and the same architecture as Model 1, improving accuracy to 93.3% for $p_{\mathrm{T}}>2$ GeV, while remaining compact enough for hardware implementation. Model 3 is a CNN that operates on 20 time-stamps of the $y$-profile and achieves the highest overall accuracy with an additional 4% gain in signal efficiency. The 3 model efficiencies versus track $p_{\mathrm{T}}$ are shown in Fig. 1(b). In spite of Model 3’s improved performance, the complexity of time-sliced data extraction poses challenges for chip design, so Model 2 was chosen for hardware implementation. The bandwidth saving using Model 2 was estimated to be in the range $57-76\%$.
@@ 3 On-chip Implementation
A smart pixel prototype ROIC was designed as a collections of 2x2 pixel analog islands surrounded by digital logic where the filtering NN sits. The ROIC was taped-out in a 1.5  mm2 ASIC chip in CMOS 28 nm bulk process. The design is shown in Figure 2(a) and a brief summary of key aspects is provided below.
@@
The NN was translated to a hardware implementation using hls4ml [9], an open-source Python framework that facilitates the co-design of ML algorithms for hardware deployment, supporting models from quantized models from QKeras and other formats [10]. We fine-tuned the numerical precision and the hardware parallelism to optimize area, performance, and power consumption. The conversion process started with the quantized model of the classifier, which hls4ml translated into HLS-ready C++ code for Siemens Catapult HLS [11] that generates a hardware description at the register-transfer level (RTL) suitable for the ASIC flow. We chose to fully parallelize the hardware logic to minimize the latency of the neural network, integrating the HLS-generated RTL design with system registers and data movers for efficient operation. A sketch of the data flow through the digital implementation is shown in Figure 2(b). The NN consumes around 300 $\mu$W. The overall power consumption per pixel, including analog and digital functions, is 6 $\mu$W, resulting in approximately 1 W/cm2, within the permissible limits of the HL-LHC experiments [12].

### v2

@@ 2 Single Cluster $p_{\mathrm{T}}$ Filtering Algorithm
Three models were developed with varying complexity to test how additional information improves $p_{\mathrm{T}}$ discrimination. The models were built to predict whether a cluster arose from three classes of charged particles: $p_{\mathrm{T}}>200$ MeV, and $p_{\mathrm{T}}<200$ MeV for positively and negatively charged particles. Two background classes were used because oppositely charged particles in a magnetic field were found to leave different cluster shapes. The threshold of 200 MeV was chosen to ensure flat signal efficiency for tracks with $p_{\mathrm{T}}>2$ GeV, a threshold important for physics analysis based on min-bias studies [7]. Model 1 is a two layer dense nueral network (DNN) with only 2 inputs , the cluster position ($y_{0}$) and size $y$-size. It correctly selects roughly 85% of tracks with $p_{T}>2$ GeV. Model 2 uses the full cluster $y$-profile and the same architecture as Model 1, improving accuracy to 93.3% for $p_{\mathrm{T}}>2$ GeV, while remaining compact enough for hardware implementation. Model 3 is a Convolutional Neural Network (CNN) that operates on 20 time-stamps of the $y$-profile and achieves the highest overall accuracy with an additional 4% gain in signal efficiency. The 3 model efficiencies versus track $p_{\mathrm{T}}$ are shown in Fig. 1(b). In spite of Model 3’s improved performance, the complexity of time-sliced data extraction poses challenges for chip design, so Model 2 was chosen for hardware implementation. The bandwidth saving using Model 2 was estimated to be in the range $57-76\%$.
@@ 3 On-chip Implementation
A smart pixel prototype ROIC was designed as a collections of 2x2 pixels consisting of analog islands surrounded by digital logic where the filtering NN sits. Each ROIC pixel is 25x25 $\mu$m2. Two $832x$ 8 arrays of pixels were taped-out on a 1.6  mm2 ASIC chip. A TSMC CMOS 28 nm bulk process was used. The design is shown in Figure 2(a) and a brief summary of key aspects is provided below.
@@
The NN was translated to a hardware implementation using hls4ml [9], an open-source Python framework that facilitates the co-design of ML algorithms for hardware deployment, supporting models from quantized models from QKeras and other formats [10]. We fine-tuned the numerical precision and the hardware parallelism to optimize area, performance, and power consumption. The conversion process started with the quantized model of the classifier, which hls4ml translated into HLS-ready C++ code for Siemens Catapult HLS [11] that generates a hardware description at the register-transfer level (RTL) suitable for the ASIC flow. We chose to fully parallelize the hardware logic to minimize the latency of the neural network, integrating the HLS-generated RTL design with system registers and data movers for efficient operation. A sketch of the data flow through the digital implementation is shown in Figure 2(b). The NN consumes around 300 $\mu$W per $832x$ 8 pixel array, assuming an estimated occupancy of 1 hit per mm2. The overall power consumption per pixel, including analog and digital functions, is 6 $\mu$W, resulting in approximately 1 W/cm2, within the permissible limits of the HL-LHC experiments [12].